OpenStudy (anonymous):

Please HElP.. The product of a number and the number increased by 9 is 15. What is the larger of the 2 solutions?

OpenStudy (anonymous):

let the number be x then the number increased by 9 is (x+9) and since their product is 15, x*(x+9) =15 solve for x and get the larger value

OpenStudy (anonymous):

do i distrubute that ??? @sauravshakya

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Yes distribute the 'x' into the parenthesis

OpenStudy (anonymous):

so its 9x ??

OpenStudy (johnweldon1993):

x times x = ? x times 9 = ?

OpenStudy (anonymous):

0

OpenStudy (johnweldon1993):

No....when you distribute the 'x' into the parenthesis...you multiply the x by each term in the parenthesis.. x(x + 9) means (x times x) + (x times 9) so first...what is x times x?

OpenStudy (anonymous):

idk wouldnt be x?

OpenStudy (johnweldon1993):

Not quite... \[\large x \times x = x^2\]

OpenStudy (anonymous):

oh right.

OpenStudy (anonymous):

9x is the next one ?

OpenStudy (johnweldon1993):

Yeah...so now that and 9 times x = 9x means we have \[\large x^2 + 9x = 15\] now we have a quadratic...can you solve this?

OpenStudy (anonymous):

do i solve itl ike that or use a formula ??..

OpenStudy (johnweldon1993):

You can either use the "completing the square method" or use the quadratic formula...which would you like to do?

OpenStudy (anonymous):

quadratic formula

OpenStudy (johnweldon1993):

Good because the completing the square would have gotten messy quick lol Do you know the quadratic formula??

OpenStudy (anonymous):

"ax^2 + bx + c = 0"

OpenStudy (anonymous):

is this it

OpenStudy (johnweldon1993):

That is the form you would use....so good lets start with that...your equation is \[\large x^2 + 9x = 15\] How do you make that look like \[\large ax^2 + bx + c = 0\] ?

OpenStudy (anonymous):

x+9=15 ?

OpenStudy (johnweldon1993):

Not quite...the only difference between your equation and the form we need is that the form we need is = to 0....yours is = to 15... so to change your equation ...into one that is = to 0....we just subtract 15 from both sides of the equation \[\large x^2+ 9x - 15 = 15 - 15\] simplified down it is... \[\large x^2+ 9x - 15 = 0\] Now...compare that to the form that we wanted... \[\large ax^2 + bx + c\] What are you 'a' 'b' and 'c' values?

OpenStudy (anonymous):

a=x b=9 c=-15

OpenStudy (johnweldon1993):

ohhh so close...a = 1...remember those values are the coefficients of the variables....so ax² ....the a = 1...because it is just 1x² right? so a = 1, b = 9, and c = -15 Okay?

OpenStudy (johnweldon1993):

NOW...we use the quadratic formula \[\huge \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] Just substitute the values you have for 'a' 'b' and 'c' into this equation...

OpenStudy (anonymous):

that sign next to -b how do you put that in the calc ? and is x always 1 ??

OpenStudy (johnweldon1993):

Dont worry about that sign for the moment...you'll see what that does soon... I'll put all the values in for you and show you... \[\huge \frac{ -9 \pm \sqrt{9^2 - 4(1)(-15)} }{ 2(1) }\] Now...we simplify...

OpenStudy (johnweldon1993):

\[\huge \frac{ -9 \pm \sqrt{81 + 60} }{ 2 }\]

OpenStudy (johnweldon1993):

This now becomes \[\huge \frac{ -9 \pm \sqrt{141} }{ 2 }\]

OpenStudy (johnweldon1993):

What you do now...is you seperate this into 2 equations...this is what that +/- sign does... This becomes... \[\huge \frac{ -9 + \sqrt{141} }{ 2 } = \space ?\] and \[\huge \frac{ -9 - \sqrt{141} }{ 2 } = \space ?\]

OpenStudy (anonymous):

-3.062828956 for the first oneee ??

OpenStudy (anonymous):

and -14.93717104 for the 2nd..

OpenStudy (johnweldon1993):

hmm...not quite what I get... \[\huge \sqrt{141} = 11.8743420870\] so \[\huge \frac{ -9 + 11.8743420870 }{ 2 } = ?\] and \[\huge \frac{ -9 = 11.8743420870 }{ 2 } = ?\]

OpenStudy (johnweldon1993):

Argh typo there...the '=' in the second is supposed to be a '-'

OpenStudy (johnweldon1993):

So I get 1.4371710435 for the first one... and -10.437171043 for the second one...so which is the largest of the 2?

OpenStudy (anonymous):

the first one thats positive ??

OpenStudy (johnweldon1993):

Mmhmm :) that would be correct!

OpenStudy (anonymous):

ok but i don't think thats the final abswer that long decimal ??

OpenStudy (anonymous):

so should i just put 1 ??

OpenStudy (johnweldon1993):

You're most likely right lol...but no I wouldn't round it down to 1...I would keep maybe 2 decimal places 1.44 I would say ...but just so we're clear...this is an approximate answer If we could I would have kept it at \[\large \frac{ -9 + \sqrt{141} }{ 2 }\] But if we want a decimal..then yeah do 1.44

OpenStudy (anonymous):

Its wrong :(

OpenStudy (johnweldon1993):

Well it is right...just must be because we rounded....because if you plug in the actual thing we got...the (-9 + √141)/2 into your equation...you get 15 maybe you should put the full decimal then... 1.437171043518959

OpenStudy (anonymous):

it said this when i got it wrong Translate to an equation. Get the equation equal to 0 before solving.

OpenStudy (johnweldon1993):

Right...we did that...remember...we make your equation = 0 and we got all the way to here...

OpenStudy (anonymous):

yes i remember that ... but why is it wrong. /:

OpenStudy (johnweldon1993):

we have the right answer...it must just be because we rounded oddly....it should be either that final equation we got... \[\large \frac{ -9 + \sqrt{141} }{ 2 }\] or the decimal equivalency of that...1.437171043518959

OpenStudy (anonymous):

did we translate this into an equation ? and trust me it will mark it wrong if i put that long number there

OpenStudy (anonymous):

i only have 2 more attempts to get this right

OpenStudy (johnweldon1993):

mmhmm....translate into equation is where we went from the words to the x(x + 9) = 15 thing...

OpenStudy (johnweldon1993):

Then set the equation = to 0 is where we distributed and subtracted 15 from both sides remember and ended up with x² + 9x - 15 = 0

OpenStudy (johnweldon1993):

I'm telling you ....we have the correct answer....just don't know how the answering service wants the answer...

OpenStudy (anonymous):

Your right did we do any of this ? Short answer questions are very particular on how to enter answers so they are graded correctly. Read the question carefully to ensure you answer the question being asked. Here are some tips to increase the chance of the short answer questions that follow be graded correctly. Don't include spaces. Such as x + 3 may be marked wrong but x+3 is correct. Don't include symbols. Such as $25 is wrong but 25 is correct. 2x is correct but 2*x or 2(x) are marked wrong. If the question asks for an expression, then = is never used. If the question asks for an equation, then = must be used. Don't round unless stated to do so in the problem. If the answer is a fraction, write the fraction instead of decimal unless the problem states otherwise. Always write in descending order, unless the problem specifies otherwise.

OpenStudy (johnweldon1993):

"If the answer is a fraction, write the fraction instead of decimal unless the problem states otherwise" This leads me to believe they want something like what we got... \[\large \frac{ -9 + \sqrt{141} }{ 2 }\]

OpenStudy (anonymous):

I think they would want just a whole number

OpenStudy (johnweldon1993):

"Don't round unless stated to do so in the problem. " I don't think so..ugh what an annoying service lol

OpenStudy (anonymous):

tell me about it had to deal with this forever lol

OpenStudy (johnweldon1993):

Haha...idk I'm thinking it wants the fraction somehow idk lol it's annoying as anything but...

OpenStudy (anonymous):

i took a practice test on this and there were no fractions!!! They were either whole numbers or decimals like this 1.6

OpenStudy (anonymous):

should i try 1.43

OpenStudy (johnweldon1993):

Lol Well idk...if you want to try 1 like you were saying...then good luck xD haha I don't know what they want... and no definietly not....1.437 does not go to 1.43....argh this is terrible....

OpenStudy (anonymous):

why wouldn't 1.43 be right & its okay I'll figure it out.

OpenStudy (johnweldon1993):

Well because...the full number is 1.437.....etc.... you would be rounding up....you cannot just cut off after the 3....the 7 DOES have an effect on that 3....it would make it round up......I know the question says don't round but....if that were the case then 1.4 would be correct to....idk

OpenStudy (anonymous):

maybe it is 1.4